# Forecasting Fold Bifurcations through Physics-Informed Convolutional Neural Networks

### v1

@@ 3.2 Mass-on-moving-belt
Figure 3: Mechanical model of the mass-on-moving-belt system (Eq. (2)).
We consider the classical mass-on-moving-belt system illustrated in Fig. 3. This system is an archetypal model for friction-induced vibrations papangelo2017subcritical ; hu2020friction . It is typically used for studying violin string dynamics leine1998stick and brake squeal kinkaid2003automotive .
The non-dimensional equation of motion of the system is as follows hu2020friction :
@@
where $v_{\text{rel}}=v-\dot{x}$ is the relative velocity. The friction force presents a discontinuity for $v_{\text{rel}}=0$. The adopted parameter values are $\mu_{\text{s}}=1$, $\mu_{\text{d}}=0.5$, $v_{0}=0.5$ and $\zeta=0.05$, which are the same parameter values utilized in papangelo2017subcritical ; hu2020friction ; habib2023predicting .
For any belt velocity, the system presents an equilibrium for which the spring and the friction force are equal and opposite. However, for belt velocities $v$ larger than a specific $v_{\text{cr}}$, the equilibrium is stable, while for $v<v_{\text{cr}}$ it is unstable. The instability is marked by a subcritical Andronov-Hopf bifurcation papangelo2017subcritical . At the Andronov-Hopf bifurcation, a branch of unstable periodic solution arises, which then merges with a branch of stable periodic solutions through a saddle-node bifurcation, as illustrated in Figure 1(b). The velocity corresponding to the saddle-node bifurcation is denoted here with $v^{*}$. For the adopted parameter values, $v_{\text{cr}}=1.151$ and $v^{*}=1.83$. The branches of stable and unstable periodic solutions merge in a non-smooth way because of the system’s discontinuity. The stable periodic solutions are characterized by a stick-slip motion.
@@ 3.3 Van der Pol-Duffing oscillator with an attached dynamic vibration absorber
The third model we consider is a van der Pol-Duffing oscillator with an attached dynamic vibration absorber. This system was thoroughly studied in the literature gattulli2001simple ; gattulli2003one , since the van der Pol oscillator is an archetypal system for self-excited oscillations candido2020non , while the Duffing term provides a nonlinearity in the stiffness, which makes the system natural frequency amplitude-dependent kovacic2011duffing . Accordingly, several studies investigated the suppression and mitigation of these kinds of self-excited oscillations through dynamic vibration absorbers habib2015suppression .
@@
$x_{1}$ and $x_{2}$ are the state variables, respectively marking the displacement of the van der Pol-Duffing oscillator and of the vibration absorber, $r$ is the mass ratio, $\gamma$ is the natural frequency ratio, $\mu_{1}$ is the van der Pol negative damping coefficient, $\mu_{2}$ is the absorber’s pseudo damping ratio and $\alpha$ the nonlinear restoring force coefficient. For a better description of the system and the derivation of the equations of motion, we address the interested reader to habib2015suppression .
@@ 3.4 Pitch-and-plunge wing profile undergoing flutter instability
The fourth system considered is a pitch-and-plunge wing profile exposed to an airflow and undergoing flutter oscillations; the system is similar to the one studied in malher2017flutter . The pitch-and-plunge wing model considered was implemented in various previous studies dowell2014modern ; lee2007suppressing1 ; lee2007suppressing2 with only slight variations from the one considered here, regarding the system nonlinearities. Non-dimensional equations of motion governing the dynamics of the system are
@@
$\begin{split}&\mathbf{x}=\left[\begin{array}[]{c}y\\ \alpha\end{array}\right],\,\mathbf{K}=\left[\begin{array}[]{ccc}\Omega^{2}&\beta u^{2}\\ 0&r_{\alpha}^{2}-\nu u^{2}\end{array}\right],\,\mathbf{C}=\left[\begin{array}[]{ccc}\zeta_{h}+\beta u0\\ -\nu u&\zeta_{\alpha}\end{array}\right],\\ &\mathbf{M}=\left[\begin{array}[]{ccc}1&x_{\alpha}\\ x_{\alpha}&r_{\alpha}^{2}\end{array}\right],\,\mathbf{b}=\left[\begin{array}[]{c}0\\ \xi_{\alpha 3}\alpha^{3}+\xi_{\alpha 5}\alpha^{5}\end{array}\right],\end{split}$ (8)
$\alpha$ marks the pitch rotation, and $y$ indicates the heave displacement, non-dimensionalized in relation to the semichord of the airfoil, while $u$ is the non-dimensional flow velocity. For the physical meaning of all the other parameters and a schematic mechanical model, we address the interested reader to malher2017flutter ; a similar nomenclature is utilized here for facilitating the comparison. The adopted parameter values are $x_{\alpha}=0.2$, $r_{\alpha}=0.5$, $\beta=0.2$, $\nu=0.08$, $\Omega=0.5$, $\zeta_{\alpha}=0.01$, $\zeta_{h}=0.01$, $\xi_{\alpha 3}=-1$ and $\xi_{\alpha 5}=20$.
@@ 4 Methods
CNNs are a particular type of machine learning architecture that has the ability to automatically extract relevant features and patterns from raw data. This ability is obtained through a series of convolutional layers that function as feature detectors. They were first introduced in the 1980s and 1990s by several researchers in the field of computer vision. One of the earliest forms of CNN is the Neocognitron, developed by Fukushima in 1980 fukushima1980neocognitron , which was inspired by the structure and function of the visual cortex and was designed to recognize patterns in visual data. Another important milestone towards the development of modern CNN is represented by the LeNet-5 architecture, developed by LeCun et al. in 1998 lecun1998gradient , which was initially used for check recognition systems for the banking industry. Despite these developments, CNNs did not become widely adopted until the 2010s, mainly because of the lack of efficient training algorithms, regularization techniques, and the availability of larger datasets. A breakthrough in CNNs was marked by the AlexNet CNN, developed by Krizhevsky et al. krizhevsky2012imagenet in 2012, which achieved a significant improvement in image classification accuracy and marked a turning point in the popularity and adoption of CNNs li2021survey .
Today, CNNs are used in a wide range of applications, including image classification, object detection, image segmentation, natural language processing, video processing, and speech recognition. They are an obvious choice for time series classification tasks.
@@ 4.1 Convolutional neural network architecture
    A fully connected layer with neurons equal to the number of classes is employed for classification.
@@
The network was trained using the Adam optimization algorithm with a mini-batch size of 32. Training progressed over a maximum of 50 epochs. The architecture of the CNN is represented in Fig. 4.
@@
Hyperparameters were tuned in a trial-and-error manner. However, the study does not aim to find the best possible network architecture but rather to analyze the importance of preprocessing the data through physically based assumptions, as explained in Sect. 4.2. Indeed, the same network architecture was used for all types of data normalization tested.
@@ 4.2 Physical information through input manipulation
We implemented several normalization approaches to investigate the importance of providing data in the correct form. This is an essential step for providing physical insight to the CNN and enabling extrapolation capabilities. In a recent work habib2023predicting , we illustrated that time series close to a fold bifurcations dissipate energy in a particular way, which is not logarithmic as for linear systems. Conversely, energy dissipation slows down in the vicinity of the fold. This phenomenon can be understood comparing Figs. 4(a) and 4(b). By depicting the logarithmic decrement of the oscillation amplitudes and measuring its minimal value during the decay, it is possible to predict the fold bifurcation, as described in habib2023predicting . According to this procedure, the oscillation frequency can be disregarded, while only the oscillation amplitude trend is relevant. Besides, amplitude represented in a logarithmic scale better visualizes the phenomenon occurring near the fold.
@@
Figure 5: Various normalization of the signal. Left column: far from the fold, center column: close to the fold, right column: after the fold. Rows indicate, in order, the original signal, min-max normalization, transformation into polar form and normalization, transformation into polar form and logarithmic scale after normalization. Trajectories were obtained for the system in (1), with $c_{1}=0.1$, $c_{3}=0.0889$, $c_{3}=0.4222$ and $c_{3}=0.4889$; $c_{3\text{cr}}=0.444$.
@@
    Min-max normalization: the signal is scaled such that the largest value is 1 and the smallest one is -1, as illustrated in Figs. 4(d)-4(f). This normalization does not provide any physical insight to the CNN, but it helps deal with systems with very different amplitude ranges. It is a standard way of normalizing data and is unrelated to any physical insight of the phenomenon.
@@
    Polar transformation and normalization: system coordinates are transformed into polar form according to the formula $\rho=\sqrt{x^{2}+\dot{x}^{2}}$, then $\rho$ is scaled such that its maximum value is 1. The $\rho$ signal provided to the network has almost no information about the system frequency as it is not an oscillatory signal anymore. This can be beneficial as frequency is not strictly related to the vicinity of the fold bifurcation. Figures 4(g)-4(i) illustrate the transformed signals. Far from the fold (Fig. 4(g)), the signal decreases regularly until zero; close to the fold (Fig. 4(h)), it presents a sort of bump before settling to 0; after the fold (Fig. 4(i)) it rapidly converges to a constant value different from zero.
@@
    Polar transformation, normalization and logarithmic scale: system coordinates are transformed in polar form according to the formula $\rho=\sqrt{x^{2}+\dot{x}^{2}}$, then $\rho$ is scaled such that its maximum value is 1, and finally it is transformed to a natural logarithmic scale, as illustrated in Figs. 4(j)-4(l). This normalization is similar to the previous one (without moving mean), with the additional feature that far from the bifurcation the signal is almost reduced to a straight line (it would be perfectly straight in the case of a linear single-DoF system). Besides, the signal does not converge to zero but to a negative number because of the logarithmic scale.
@@ 5 Results
The same network was trained with the same training data for all types of normalization. In all cases, 100 % accuracy was obtained during training and validation. This result suggests that the CNN utilized and the amount of data is adequate for the task. Conversely, providing correct classification for other parameter ranges is significantly more challenging. Different tests were defined, namely:
@@
For the first task, results can be quantified with a percentage of accuracy, while only qualitative assessments are given for the other tasks. Results are summarized in Table 1, while their comprehensive discussion is provided below. In the table, the following qualitative labels are provided. If far and after trajectories are significantly wrongly classified, the network is judged as “wrong”; if close trajectories are misclassified, then the label “no close” is given, if the network is overall correct, but some errors are present it is marked as “inaccurate”; while if the network provides correct classification, it is marked as “good”.
Min-max 66.6 % wrong wrong no close
Polar 87.5 % no close good good
Pol-MovMean 87.33 % no close good good
Pol-log 100 % wrong good inaccurate
Pol-log-MovMean 100 % good good good
@@ 5.1 Min-max normalization
Figure 6: Prediction diagrams obtained by the CNN trained with min-max normalized data. Color interpretation: green-far, red-close, blue-after.
@@ 5.2 Polar transformation and normalization
Figure 7: Prediction diagrams obtained by the CNN trained with data normalized in polar form. Color interpretation: green-far, red-close, blue-after.
The polar form transformation and amplitude normalization between 0 and 1 provided much better results. This is mostly due to the almost complete elimination of the frequency of oscillation from the data, which is not related to general traits of system approaching a fold bifurcation. We also note that the network never misclassified far and after trajectories. Regarding the system with nonlinear damping (Fig. 6(a)), reducing $c_{1}$ the network significantly reduced the trajectories classified as close, which are limited to those extremely close to the fold. About the mass-on-moving-belt system (Fig. 6(b)), the network completely overlooked close trajectories, which were always classified as far. Regarding the van der Pol-Duffing oscillator (Fig. 6(c)) and the pitch-and-plunge wing profile (Fig. 6(d)) the classification was excellent with both degrees of freedom.

### v2

@@ 3.2 Mass-on-moving-belt
(a) Mass-on-moving-belt system (Eq. (2))
(b) Pitch-and-plunge airfoil (Eq. (7))
Figure 3: Mechanical models of the mass-on-moving-belt system (a) and pitch-and-plunge airfoil (b).
We consider the classical mass-on-moving-belt system illustrated in Fig. 3(a). This system is an archetypal model for friction-induced vibrations [11, 12]. It is typically used for studying violin string dynamics [74] and brake squeal [75].
The non-dimensional equation of motion of the system is as follows [12]:
@@
where $v_{\text{rel}}=v-\dot{x}$ is the relative velocity. The friction force presents a discontinuity for $v_{\text{rel}}=0$. The adopted parameter values are $\mu_{\text{s}}=1$, $\mu_{\text{d}}=0.5$, $v_{0}=0.5$ and $\zeta=0.05$, which are the same parameter values utilized in [11, 12, 45].
For any belt velocity, the system presents an equilibrium for which the spring and the friction force are equal and opposite. However, for belt velocities $v$ larger than a specific $v_{\text{cr}}$, the equilibrium is stable, while for $v<v_{\text{cr}}$ it is unstable. The instability is marked by a subcritical Andronov-Hopf bifurcation [11]. At the Andronov-Hopf bifurcation, a branch of unstable periodic solution arises, which then merges with a branch of stable periodic solutions through a saddle-node bifurcation, as illustrated in Figure 2(b). The velocity corresponding to the saddle-node bifurcation is denoted here with $v^{*}$. For the adopted parameter values, $v_{\text{cr}}=1.151$ and $v^{*}=1.83$. The branches of stable and unstable periodic solutions merge in a non-smooth way because of the system’s discontinuity. The stable periodic solutions are characterized by a stick-slip motion.
@@ 3.3 Van der Pol-Duffing oscillator with an attached dynamic vibration absorber
The third model we consider is a van der Pol-Duffing oscillator with an attached dynamic vibration absorber. This system was thoroughly studied in the literature [76, 77], since the van der Pol oscillator is an archetypal system for self-excited oscillations [78], while the Duffing term provides a nonlinearity in the stiffness, which makes the system natural frequency amplitude-dependent [79]. Accordingly, several studies investigated the suppression and mitigation of these kinds of self-excited oscillations through dynamic vibration absorbers [80].
@@
$x_{1}$ and $x_{2}$ are the state variables, respectively marking the displacement of the van der Pol-Duffing oscillator and of the vibration absorber, $r$ is the mass ratio, $\gamma$ is the natural frequency ratio, $\mu_{1}$ is the van der Pol negative damping coefficient, $\mu_{2}$ is the absorber’s pseudo damping ratio and $\alpha$ the nonlinear restoring force coefficient. For a better description of the system and the derivation of the equations of motion, we address the interested reader to [80].
@@ 3.4 Pitch-and-plunge wing profile undergoing flutter instability
The fourth system considered is a pitch-and-plunge wing profile exposed to an airflow and undergoing flutter oscillations (Fig. 3(b)); the system is similar to the one studied in [81]. The pitch-and-plunge wing model considered was implemented in various previous studies [82, 83, 84] with only slight variations from the one considered here, regarding the system nonlinearities. Non-dimensional equations of motion governing the dynamics of the system are
@@
$\begin{split}&\mathbf{x}=\left[\begin{array}[]{c}y\\ \alpha\end{array}\right],\,\mathbf{K}=\left[\begin{array}[]{ccc}\Omega^{2}&\beta u^{2}\\ 0&r_{\alpha}^{2}-\nu u^{2}\end{array}\right],\,\mathbf{C}=\left[\begin{array}[]{ccc}\zeta_{h}+\beta u&0\\ -\nu u&\zeta_{\alpha}\end{array}\right],\\ &\mathbf{M}=\left[\begin{array}[]{ccc}1&x_{\alpha}\\ x_{\alpha}&r_{\alpha}^{2}\end{array}\right],\,\mathbf{b}=\left[\begin{array}[]{c}0\\ \xi_{\alpha 3}\alpha^{3}+\xi_{\alpha 5}\alpha^{5}\end{array}\right],\end{split}$ (8)
$\alpha$ marks the pitch rotation, and $y$ indicates the heave displacement, non-dimensionalized in relation to the semichord of the airfoil, while $u$ is the non-dimensional flow velocity. For the physical meaning of all the other parameters, we address the interested reader to [81]; a similar nomenclature is utilized here for facilitating the comparison. The adopted parameter values are $x_{\alpha}=0.2$, $r_{\alpha}=0.5$, $\beta=0.2$, $\nu=0.08$, $\Omega=0.5$, $\zeta_{\alpha}=0.01$, $\zeta_{h}=0.01$, $\xi_{\alpha 3}=-1$ and $\xi_{\alpha 5}=20$.
@@ 4 Methods
CNNs are a particular type of machine learning architecture that has the ability to automatically extract relevant features and patterns from raw data. This ability is obtained through a series of convolutional layers that function as feature detectors. They were first introduced in the 1980s and 1990s by several researchers in the field of computer vision. One of the earliest forms of CNN is the Neocognitron, developed by Fukushima in 1980 [85], which was inspired by the structure and function of the visual cortex and was designed to recognize patterns in visual data. Another important milestone towards the development of modern CNN is represented by the LeNet-5 architecture, developed by LeCun et al. in 1998 [86], which was initially used for check recognition systems for the banking industry. Despite these developments, CNNs did not become widely adopted until the 2010s, mainly because of the lack of efficient training algorithms, regularization techniques, and the availability of larger datasets. A breakthrough in CNNs was marked by the AlexNet CNN, developed by Krizhevsky et al. [87] in 2012, which achieved a significant improvement in image classification accuracy and marked a turning point in the popularity and adoption of CNNs [56].
Today, CNNs are used in a wide range of applications, including image classification, object detection, image segmentation, natural language processing, video processing, and speech recognition. They are an optimal solution for time series classification tasks.
@@ 4.1 Convolutional neural network architecture
    A fully connected layer with an input size of 64 and output of 3 is employed for classification.
@@
The network was trained using the Adam optimization algorithm with a mini-batch size of 32. Training progressed over a maximum of 50 epochs. Initial learning rate is set to 0.001. Padding direction is set to ‘left’; however, all input signals completely filled the required signal length. The architecture of the CNN is represented in Fig. 4.
@@
Hyperparameters were tuned in a trial-and-error manner and no analysis of the network sensitivity to the hyperparameters was performed. However, the study does not aim to find the best possible network architecture but rather to analyze the importance of pre-processing the data through physically based assumptions, as explained in Sect. 4.2. Indeed, the same network architecture was used for all types of data normalization tested. All computation for this research was computed in the MATLAB environment.
@@ 4.2 Physical information through input manipulation
We implemented several normalization approaches to investigate the importance of providing data in the correct form. This is an essential step for providing physical insight to the CNN and enabling extrapolation capabilities. In a recent work [45], we illustrated that time series close to a fold bifurcations dissipate energy in a particular way, which is not logarithmic as for linear systems. Conversely, energy dissipation slows down in the vicinity of the fold. This phenomenon can be understood comparing Figs. 5(a) and 5(b). By depicting the logarithmic decrement of the oscillation amplitudes and measuring its minimal value during the decay, it is possible to predict the fold bifurcation, as described in [45]. According to this procedure, the oscillation frequency can be disregarded, while only the oscillation amplitude trend is relevant. Besides, amplitude represented in a logarithmic scale better visualizes the phenomenon occurring near the fold.
@@
Figure 5: Various normalization of the signal. Left column: far from the fold, center column: close to the fold, right column: after the fold. Rows indicate, in order, the original signal, min-max normalization, transformation into polar form and normalization, transformation into polar form and logarithmic scale after normalization. Trajectories were obtained for the system in (1), with $c_{1}=0.1$, $c_{3}=0.0889$, $c_{3}=0.4222$ and $c_{3}=0.4889$; $c_{3\text{cr}}=0.444$. The length of the time series in the figure does not strictly reflect the length of the time series used for the training.
@@
    Min-max normalization: the signal is scaled such that the largest value is 1 and the smallest one is -1, as illustrated in Figs. 5(d)-5(f). This normalization does not provide any physical insight to the CNN, but it helps deal with systems with very different amplitude ranges. It is a standard way of normalizing data and is unrelated to any physical insight of the phenomenon.
@@
    Polar transformation and normalization: system coordinates are transformed into polar form according to the formula $\rho=\sqrt{x^{2}+\dot{x}^{2}}$, then $\rho$ is scaled such that its maximum value is 1. The $\rho$ signal provided to the network has a reduced information about the system frequency. This can be beneficial as frequency is not strictly related to the vicinity of the fold bifurcation. Figures 5(g)-5(i) illustrate the transformed signals. Far from the fold (Fig. 5(g)), the signal decreases regularly until zero; close to the fold (Fig. 5(h)), it presents a sort of bump before settling to 0; after the fold (Fig. 5(i)) it rapidly converges to a constant value different from zero.
@@
    Polar transformation, normalization and logarithmic scale: system coordinates are transformed in polar form according to the formula $\rho=\sqrt{x^{2}+\dot{x}^{2}}$, then $\rho$ is scaled such that its maximum value is 1, and finally it is transformed to a natural logarithmic scale, as illustrated in Figs. 5(j)-5(l). This normalization is similar to the previous one (without moving mean), with the additional feature that far from the bifurcation the signal is almost reduced to a straight line (it would be perfectly straight in the case of a linear single-DoF system). Besides, the signal does not converge to zero but to a negative number because of the logarithmic scale.
@@ 4.3 Training
Figure 6: Graphical representation of the parameter values utilized for training and validation. Vertical lines indicate boundary of regions, dots mark selected values for training and validation. Green: far, red: close, blue: after.
We remark that the lack of normalization regarding the signal length makes it practically impossible for the NN to identify the signal’s frequency. This choice reflects the practical difficulty in real applications of selecting signals with the same time length normalized with respect to the system’s frequency. Nevertheless, as discussed in Sect. 7, the signal frequency has almost no correlation with the classification proposed for the time series (far, close or after the fold). In fact, for all normalization proposed, except min-max, the information about the system frequency is purposely removed. The only normalization that might be significantly affected by the lack of signal length normalization is the min-max; however, performing the same analysis discussed below with time-length normalized signal did not improve the min-max normalized network performance. The analysis is omitted here for the sake of brevity.
@@ 5 Results
Networks were trained with the same training data for all types of normalization. In all cases, 100 % accuracy was obtained during training and validation. This result suggests that the NN architecture and the amount of data are adequate for the task. Indeed, 100 % accuracy is quite unusual for any machine learning problem. In the authors’ opinion, two main factors lead to this result. First of all, the validation set is relatively small, encompassing only 600 samples. However, considering that the training data included only 3000 time series, it seems appropriate. The second reason is that the trajectories are very similar to each other within each class, and very different between classes. Most probably, reducing the gap between far and close regions would reduce the accuracy of the NNs validation. Figure 6 illustrates that the validation data cover the whole $c_{3}$ range of each class.
Min-max 66.6 % wrong wrong no close
Polar 87.5 % no close good good
Pol-MovMean 87.33 % no close good good
Pol-log 100 % wrong good inaccurate
Pol-log-MovMean 100 % good good good
Providing correct classification for other parameter ranges is significantly more challenging. Different tests were defined, namely:
@@
For the first task, results can be quantified with a percentage of accuracy, while only qualitative assessments are given for the other tasks because of the arbitrarity of the boundary between far and close regions. Results are summarized in Table 1, while their comprehensive discussion is provided below. In the table, the following qualitative labels are provided. If far and after trajectories are significantly wrongly classified, the network is judged as “wrong”; if close trajectories are misclassified, then the label “no close” is given, if the network is overall correct, but some errors are present it is marked as “inaccurate”; while if the network provides correct classification, it is marked as “good”.
Type of normalization Mass-on mov.-belt VdP Duf. $x_{1}$ VdP Duf. $x_{2}$ Pitch & plunge $y$ Pitch & plunge $\alpha$
Min-max - - - - -
Polar - 0.234 0.359 0.690 0.234
Pol-MovMean - 0.199 0.352 0.695 0.227
Pol-log - 0.504 0.640 - -
Pol-log-MovMean 0.368 0.552 0.577 0.965 0.948
Pol-log-MovMean (noisy test data) - 0.993 0.804 0.846 0.857
Pol-log-MovMean (noisy test and training data) 0.722 - 0.497 0.447 0.169
Table 2: Ratio between the overlapping area of estimated close-far regions and close region.
@@
Since the boundary between far and close regions is not well defined, and the shape of the time series depends on their initial conditions, the NN provides a region of overlapping between far and close region, resulting in a fuzzy boundary. In order to quantify the overlapping, the ratio between the overlapping region and the full extent of the estimated close region is calculated, as indicated in Table 2 for all types of normalization implemented. The smaller this number, the better the boundary between the two regions is defined. Interestingly, we note that, although the final pre-processing strategy (transformation into polar coordinates, normalization, transformation into a logarithmic scale, and filtering via a moving average) was the only one able to correctly classify trajectories from all systems, it resulted in a much larger overlapping region, significantly larger than when the logarithmic transformation was not applied. This is clearly an undesirable outcome. As a result, a clear best normalization strategy cannot be defined, as discussed in Sect. 7.
@@ 5.1 Min-max normalization
Figure 7: Prediction diagrams obtained by the CNN trained with min-max normalized data. Each point corresponds to a different time series. Dashed lines indicate the fold bifurcation. The colors mark the predicted classification. Color interpretation: green-far, red-close, blue-after.
@@ 5.2 Polar transformation and normalization
Figure 8: Prediction diagrams obtained by the CNN trained with data normalized in polar form. Each point corresponds to a different time series. Dashed lines indicate the fold bifurcation. The colors mark the predicted classification. Color interpretation: green-far, red-close, blue-after.
The polar form transformation and amplitude normalization between 0 and 1 provided much better results. This is mostly due to the almost complete elimination of the frequency of oscillation from the data, which is not related to general traits of system approaching a fold bifurcation. We also note that the network never misclassified far and after trajectories. Regarding the system with nonlinear damping (Fig. 8(a)), reducing $c_{1}$ the network significantly reduced the trajectories classified as close, which are limited to those extremely close to the fold. About the mass-on-moving-belt system (Fig. 8(b)), the network completely overlooked close trajectories, which were always classified as far. Regarding the van der Pol-Duffing oscillator (Fig. 8(c)) and the pitch-and-plunge wing profile (Fig. 8(d)) the classification was excellent with both degrees of freedom. The overlapping region between far and close is also relatively small, except for the pitch-and-plunge if computed from $y$ (Table 2).